### v1

# A SSM is polymerized from multivariate time series

Abstract
For multivariate time series (MTS) tasks, previous state space models (SSMs) followed the modeling paradigm of Transformer-based methods. However, none of them explicitly model the complex dependencies of MTS: the Channel Dependency variations with Time (CDT). In view of this, we delve into the derivation of SSM, which involves approximating continuously updated functions by orthogonal function bases. We then develop Poly-Mamba, a novel method for MTS forecasting. Its core concept is to expand the original orthogonal function basis space into a multivariate orthogonal function space containing variable mixing terms, and make a projection on this space so as to explicitly describe the CDT by weighted coefficients. In Poly-Mamba, we propose the Multivariate Orthogonal Polynomial Approximation (MOPA) as a simplified implementation of this concept. For the simple linear relationship between channels, we propose Linear Channel Mixing (LCM) and generate CDT patterns adaptively for different channels through a proposed Order Combining method. Experiments on six real-world datasets demonstrate that Poly-Mamba outperforms the SOTA methods, especially when dealing with datasets having a large number of channels and complex correlations. The codes and log files will be released at: https://github.com/Joeland4/Poly-Mamba.
1 Introduction
Structured state space models (SSMs) [1] are subquadratic-time foundational architectures compared with Transformers [2], and show great performance with approximately linear complexity on long-range dependency tasks. Previous studies [3] [4] [5] attempted to employ SSM for Multivariate Time Series Forecasting (MTSF), they all follow the Transformer-based MTSF modeling paradigm: learning dependencies between temporal tokens [6] [7] [8] [9], Channel tokens [10] and their concatenation [11]. However, the special complex dependency pattern of MTS is the Channel Dependency variations with Time (CDT), none of these methods explicitly depict it.
It is inappropriate to directly model the CDT because it not only greatly increases complexity when calculating the dependency between temporal tokens of all channels but is also hard to generalize for the scale of most MTS data. We delved deep into the initial development of SSM [12]: real-time approximation of continuously updating function by orthogonal function basis [13], and we found that compared with Transformers, SSM has the potential to efficiently and effectively depict the CDT pattern.
We propose an improved SSM for MTSF tasks named Poly-Mamba. Based on the dynamical equation of its fitting coefficient, we propose three operations, Multivariate Orthogonal Polynomial Approximation (MOPA), Linear Channel Mixing (LCM) and Order Combining, improving the ability of capturing the complex CDT pattern.
The idea of MOPA is to explicitly describe the inter-channel dependency pattern by using the mapping weighted coefficients of the extended multivariate polynomial function basis, which contains variable mixing terms. MOPA is an efficient operation for mapping on the multivariate polynomial basis space after simplification. LCM is a mapping operation used to construct a simple linear correlation between channels. In addition, in order to deal with different MTS scenarios, we propose Order Combining, which retains the low-order trend information of each channel itself and uses gating [14] to adaptively generate the CDT pattern of each channel.
The architecture of Poly-Mamba for MTSF takes time series patches as tokens and generates corresponding multivariate hidden states. For the multivariate hidden states at each moment, LCM, MOPA, and Order Combining are applied to transform into hidden states with increased inter-channel correlations. Thus, this new SSM process generates a pattern representation of the time-varying dependencies between multi-channel tokens.
Our contributions are as follows:
  • We delve into the derivation of SSM, improve it for MTSF, and propose a new model named Poly-Mamba, which explicitly models channel dependency variations with time.
  • We develop MOPA, an efficient method based on the idea of extending the orthogonal polynomial function basis space to a multivariate orthogonal polynomial basis space and explicitly characterizing channel dependencies through mapping weights.
  • We propose LCM to handle the simple linear correlation between channels. Additionally, we propose Order Combining, retaining the trend information of each channel itself and adaptively learning different types of CDT patterns.
  • We conduct extensive experiments on six real-world datasets. The experimental results show that our Poly-Mamba outperforms the SOTA methods for the MTSF task, proving that the improved SSM can effectively capture the complex dependencies of MTS.
2 Preliminary
In Section 2.1, we first briefly review the derivation of SSM dynamics. Subsequently, in Section 2.2, we provide a description of multivariate Legendre polynomials.
2.1 SSM
The essence of the state space model is to regard sequence pattern learning as an online function estimation problem. The form basis of its dynamic equation was first proposed in HIPPO [12]. It produces operators projecting arbitrary functions onto the space of orthogonal polynomials with respect to a given measure (Equation 1), The closed-form ODE or linear recurrence (Equation 2) allows fast incremental updating of the optimal polynomial approximation as the input function is revealed through time.
[…]
$x^{\prime}(t)=Ax(t)+Bu(t)$ (2)
$y(t)=Cx(t)+Du(t)$
The target function $u(t)$, a continuously collected signal, is approximated by a linear combination of standard orthogonal function bases $\left\{g_{n}(t)\right\}_{n=0}^{N}$. For this, HIPPO [12] uses an online function approximation method, that is, mapping each static interval $s\in[a,b]$ to the entire sequence interval $t\in[0,T]$ through the defined Measure. The analytical solution of the combination coefficient $C_{n}$ is:
[…]
Taking the HIPPO Matrix in S4 [1] corresponding to HIPPO-LegS as an example, that is, using Legendre polynomials $P_{n}$ and the corresponding standard orthogonal basis $g_{n}$, satisfying:
$\int_{-1}^{1}p_{m}(s)$ $p_{n}(s)dt=\frac{2}{2n+1}\delta_{m,n}$ (4)
$g_{n}(s)$ $=\sqrt{\frac{2n+1}{2}}p_{n}(s)$
In LegS, the measure is defined as $t_{\leq T}(s)=(s+1)T/2$. It uniformly maps the domain $[-1,1]$ of $P_{n}$ to $[0,T]$. By taking the derivative of both sides with respect to $T$ and substituting $t_{\leq T}(s)$, $P_{n}$ and $P^{\prime}_{n}$ into Equation 3 (see details in [12]), the final ODE form is obtained:
$\frac{d}{dT}c_{n}(T)=\frac{\sqrt{2(2n+1)}}{T}u(T)-\frac{1}{T}(-nc_{n}(T)+\sum_{k=0}^{n}\sqrt{(2n+1)(2k+1)}c_{k}(T))$ (5)
Putting all the coefficients $c_{n}(T)$ together and letting $x(t)=(c_{0},c_{1},\ldots,c_{N})$, the dynamic form of SSM is obtained:
$x^{\prime}(t)=\frac{A}{t}x(t)+\frac{B}{t}u(t)$ (6)
$A_{nk}=-\left\{\begin{matrix}(2n+1)^{1/2}(2k+1)^{1/2}\ if\ n>k\\ n+1\qquad\qquad\qquad\quad\ \ if\ n=k\\ 0\qquad\qquad\qquad\qquad\quad\ if\ n<k\end{matrix}\right.,$ $\ B_{n}=\sqrt{2(2n+1)}$
S4 [1] derives the HIPPO Matrix A as a diagonal matrix and gives the corresponding convolution form for efficient training. Mamba [15], introduces a selective mechanism into SSM, enabling it to discern the importance of information like the attention mechanism.
2.2 Multivariable Legendre polynomials
In the space $L_{2}(I)$, we can use univariate orthogonal polynomials to construct a set of multivariate orthogonal polynomials [16]. Specifically, consider a set of orthogonal Legendre polynomials $\varphi_{n}(x_{c})$ that are exact up to degree $N$ in a single variable $x_{c}$. Then, the set of multivariate Legendre polynomials $\varphi_{n_{1}\cdots n_{C}}(x)=\varphi_{n_{1}}(x_{1})\cdots\varphi_{n_{C}}(x_{C})$ in $C$ variables constitutes a set of multivariate orthogonal polynomials on $L_{2}(I)$. The set of multivariate polynomials $\varphi_{n_{1}\cdots n_{c}}(x)$ in $C$ variables up to degree $n_{deg}=\sum_{c=1}^{C}n_{c}$ has exactly $n_{\Phi}=\begin{pmatrix}C+n_{deg}\\ C\end{pmatrix}$ elements.
3 Poly-Mamba
Figure 1: (left) Poly-Mamba architecture. (right) Three operations on coefficients in the improved SSM.
To effectively learn the CDT pattern of MTS, Poly-Mamba models along the temporal direction. We embed sequence segment patches as tokens. This is inspired by the Transformer-based model PatchTST [9]. Treating sequence segments as tokens can be more semantically informative and can effectively reduce complexity, and this is also applicable to the Mamba. As shown in the Figure 1, in addition to the original parameters ($A$, $B$, $C$, $\Delta$) in Mamba [15], Poly-Mamba adds operations and corresponding parameters of MOPA, LCM, and Order Combining to the hidden state. The complete process in the form of recurrent is shown in Algorithm 1. For the MTSF task, after flattening the encoding result of $L$ layers, a fully connected is used to output the predictions.
3.1 MOPA
Our idea is to map the multivariate space containing multivariate mixing terms and describe the complex dependencies between channels with learnable mapping weights. The whole process consists of two steps. The first is to expand the univariate orthogonal polynomial into a multivariate orthogonal polynomial (Equation 7), and obtain the corresponding coefficients. The original basis space with $N+1$ dimensions is expanded into a basis space with $N_{\Phi}=\sum_{n_{deg=0}}^{N_{deg}}\begin{pmatrix}C+n_{deg}\\ C\end{pmatrix}$ dimensions.
$P_{n_{1},n_{2},...,n_{C}}(x_{1},x_{2},...,x_{C})=P_{n_{1}}(x_{1})P_{n_{2}}(x_{2})...P_{n_{c}}(x_{c})...P_{n_{C}}(x_{C})$ (7)
For example, when the number of channels $C=2$, a quadratic term $x_{1}^{2}$ of the original channel $x_{1}$ becomes three terms: $x_{1}^{2}$, $x_{2}^{2}$, and $x_{1}x_{2}$. Here, $x_{1}x_{2}$ represents the mixing term of two variables. The coefficients of different mixing terms represent different degrees of interdependence between channels. The second step is to apply a projection on coefficients of $N_{\Phi}$ terms:
$C_{N_{\Phi}}^{\prime}=f(Linear(C_{N_{\Phi}}))=f(W_{N_{\Phi}}C_{N_{\Phi}}+B)$ (8)
where $W_{N_{\Phi}}\in\mathbb{R}^{N_{\Phi}\times N_{\Phi}^{\prime}}$, and an activation function $f$ can be added to make it a nonlinear mapping. Although the weights in the weighted summation of the coefficients of the multivariate orthogonal polynomial function basis can intuitively describe the inter-channel dependence, the two-step calculation plus the linear layer after the dimensional power expansion will greatly increase the complexity.
MOPA makes two simplifications to approximate the above mapping process. First, the dimension $N_{\Phi}^{\prime}$ after mapping is set to be equal to $N+1$, the original dimension of the hidden space. And perform mapping for each order separately, that is, perform weighted summation on the coefficients of all terms of that order. So now the coefficients of each order in the mapped order $N$ are:
$C_{n}^{\prime}=f(\sum_{i=1}^{n_{\Phi}}w_{i}C_{n,i}+b_{i}),n\in\left\{0,1,...,N\right\}$ (9)
where $n_{\Phi}=\begin{pmatrix}C+n\\ C\end{pmatrix}$. The first simplification is due to the large amount of repetition and redundancy in previous cross-order linear relationships. Describing the interdependence between channels separately on each order is sufficient. Compared to the complete mapping of the dimension after power expansion, the complexity is significantly reduced.
But the complexity of performing projection on each order of each channel is still too high. MOPA continues to make a simplified approximation. Do not instantiate the weighted summation process and weight $w_{i}$ for each order. Instead, directly replace the expansion and projection processes with one transformation parameter. Therefore, the coefficient transformation parameter matrix $M$ of size $C\times N$ constructed by MOPA is multiplied by the Hadamard product with the hidden state (coefficient matrix) of size $C\times N$:
$MOPA(x(t))=M\cdot x(t)$ (10)
When substituted into Equation 3, the new polynomial Equation 7 and its partial derivative only add the tensor product with polynomials of other variables. Thus, they do not affect the dynamic form of SSM and the parameters in Mamba. As shown in Figure 1 right, in this way, the complex CDT pattern between channels can be concisely and efficiently modeled. Note that the transformation of direct dot product by MOPA is not a completely corresponding method but a simplified variation that can still contain the essence of the original two-step mapping process.
3.2 LCM
Different MTS scenarios exhibit varying degrees of CDT patterns. For example, some channels may exhibit a simple linear relationship. Then, for the linear relationship between channels, it is not appropriate for MOPA to capture it with refined high-order channel cross-terms. Therefore, we propose the LCM method. LCM characterizes the linear relationship between channels with a learnable parameter matrix $L$ of size $C\times C$, and multiplies $L$ with the coefficient matrix of size $C\times N$ :
$LCM(x(t))=L\ x(t)$ (11)
In this way, each channel receives information from other channels and is mapped to the polynomial space linearly related to the channels.
3.3 Order Combining
LCM and MOPA model the simple linear dependence and complex relationships between channels respectively. However, the specific type of channel relationship is unknown in advance. Order Combining adaptively generates the CDT pattern by applying a Gate operation [14] on the outputs of MOPA and LCM:
$G^{\left\{2,...,N\right\}}=softmax(P_{L}\ LCM^{\left\{2,...,N\right\}},\ P_{M}\ MOPA^{\left\{2,..,N\right\}})$ (12)
where $P_{L}$ and $P_{M}$ are the scalar weights. Moreover, in order to maintain the trend (low-order) information of each channel itself, we perform LCM mapping on all terms but only perform MOPA on high-order terms. After performing gate selection on the high-order terms of LCM and MOPA, they are spliced with the low-order terms of LCM:
$OC=Concat(LCM^{\left\{0,1\right\}},\ G^{\left\{2,...,N\right\}})$ (13)
If the overall relationship is linear, then LCM is selected by the gate; if it is a high-order complex relationship, MOPA is selected by the gate, and the coefficients of the low-order terms in front will be adjusted accordingly to adapt to the complex pattern.
Algorithm 1: The procedure of SSM in Poly-Mamba under recurrence form
$[B,C,L,D]\qquad\qquad$Input: $Batch(X)$
$[B,C,L,D]\qquad\qquad$Output: $Batch(Y)$
$\qquad\qquad\qquad\qquad\quad\ Y\leftarrow List[\ ]$
                            For $t$ in range  (1,  $L$):
$[B,C,D,N]\qquad\qquad\qquad h_{t}\ \leftarrow\ \overline{A}\ h_{t-1}+\overline{B}\ X_{t}$
$[B,D,C,N]\qquad\qquad\qquad h_{t}\leftarrow Reshape(h_{t})$
$[B,D,C,N]\qquad\qquad\qquad LCM(h_{t})\leftarrow L\ h_{t}$
$[B,D,C,N-2]\qquad\qquad\ MOPA(h_{t}[2:])\leftarrow M\cdot h_{t}$
$[B,D,C,N-2]\qquad\qquad\ MoE\leftarrow softmax(P_{L}\ast LCM(h_{t})[2:],\ P_{M}\ast MOPA(h_{t}[2:]))$
$[B,D,C,N]\qquad\qquad\qquad h^{\prime}_{t}\leftarrow Concat(LCM(h_{t})[0,1],\ MoE)$
$[B,D,C]\qquad\quad\qquad\qquad\ y_{t}\leftarrow C\ h^{\prime}_{t}$
$[B,C,D]\qquad\quad\qquad\qquad\ y_{t}\leftarrow Reshape(y_{t})$
$[B,C,L,D]\qquad\qquad\qquad Y\leftarrow Y.append(y_{t})$
                            End for $t$
4 Experiments
We comprehensively evaluate the proposed Poly-Mamba on six real-world datasets, including ETT (consisting of 4 subsets) [6], ECL [7], Exchange [7], Traffic [7], Weather [7] and Solar-Energy [17]. We select seven well-known forecasting models as our baselines, which include: (1) Mamba-based methods: S-Mamba [4]; (2) Transformer-based methods: Crossformer [11], PatchTST [9], and iTransformer [10]; (3) Linear-based methods: DLinear [18], TiDE [19]; and (4) CNN-based methods: TimesNet [20]. The experimental setting is identical to that in iTransformer [10], where the input length $L=96$ and the output length $T=\left\{96,192,336,720\right\}$.
4.1 Main Results
As is obviously shown in Table 1, by comparing the experimental results with the data visualization of different datasets, we can find that Poly-Mamba has a better prediction performance for scenarios where there is a real-time correlation between channels, such as on Weather. Of course, for scenarios where there is no strong dependence between channels, such as on ETT-m, the prediction effect is on par with that of PatchTST [9] which uses channel independence.
Secondly, one of the advantages of the state-space model is the compression of history. When facing longer prediction interval requirements, Poly-Mamba shows more stable prediction results. Moreover, compared with the S-Mamba [4] model which is also based on the state-space but models along the channel direction, Poly-Mamba shows more excellent prediction stability. For example, by comparing the results of prediction lengths $T=\left\{336,720\right\}$ on the Solar and Weather datasets, we can find that Poly-Mamba’s modeling is closer to the complex pattern of them.
Table 1: Main results for the long-term MTSF task. The experimental setting is consistent with iTransformer[10], where input sequence length $T=96$ and output sequence length $S=\left\{96,192,336,720\right\}$ for all baselines. Avg means the average results from all four prediction lengths. Red value is the best and blue with underline is the second.
Models Poly-Mamba S-Mamba iTransformer PatchTST Crossformer TiDE TimesNet DLinear
(Ours) (2024) (2024) (2023) (2023) (2023) (2023) (2023)
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.321 0.364 0.325 0.361 0.334 0.368 0.329 0.367 0.404 0.426 0.364 0.387 0.338 0.375 0.345 0.372
192 0.365 0.387 0.368 0.385 0.377 0.391 0.367 0.385 0.450 0.451 0.398 0.404 0.374 0.387 0.380 0.389
336 0.393 0.409 0.401 0.408 0.426 0.420 0.399 0.410 0.532 0.515 0.428 0.425 0.410 0.411 0.413 0.413
720 0.449 0.436 0.469 0.446 0.491 0.459 0.454 0.439 0.666 0.589 0.487 0.461 0.478 0.450 0.474 0.453
Avg 0.382 0.399 0.391 0.400 0.407 0.410 0.387 0.400 0.513 0.496 0.419 0.419 0.400 0.406 0.403 0.407
ETTm2 96 0.179 0.265 0.180 0.264 0.180 0.264 0.175 0.259 0.287 0.366 0.207 0.305 0.187 0.267 0.193 0.292
192 0.248 0.311 0.245 0.306 0.250 0.309 0.241 0.302 0.414 0.492 0.290 0.364 0.249 0.309 0.284 0.362
336 0.318 0.356 0.309 0.347 0.311 0.348 0.305 0.343 0.597 0.542 0.377 0.422 0.321 0.351 0.369 0.427
720 0.409 0.409 0.408 0.403 0.412 0.407 0.402 0.400 1.730 1.042 0.558 0.524 0.408 0.403 0.554 0.522
Avg 0.289 0.335 0.286 0.330 0.288 0.332 0.281 0.326 0.757 0.610 0.358 0.404 0.291 0.333 0.350 0.401
ETTh1 96 0.384 0.404 0.382 0.401 0.386 0.405 0.414 0.419 0.423 0.448 0.479 0.464 0.384 0.402 0.386 0.400
192 0.450 0.440 0.434 0.435 0.441 0.436 0.460 0.445 0.471 0.474 0.525 0.492 0.436 0.429 0.437 0.432
336 0.492 0.455 0.482 0.458 0.487 0.458 0.501 0.466 0.570 0.546 0.565 0.515 0.491 0.469 0.481 0.459
720 0.484 0.475 0.481 0.480 0.503 0.491 0.500 0.488 0.653 0.621 0.594 0.558 0.521 0.500 0.519 0.516
Avg 0.453 0.444 0.445 0.444 0.454 0.447 0.469 0.454 0.529 0.522 0.541 0.507 0.458 0.450 0.456 0.452
ETTh2 96 0.294 0.348 0.295 0.348 0.297 0.349 0.302 0.348 0.745 0.584 0.400 0.440 0.340 0.374 0.333 0.387
192 0.373 0.392 0.376 0.397 0.380 0.400 0.388 0.400 0.877 0.656 0.528 0.509 0.402 0.414 0.477 0.476
336 0.412 0.429 0.420 0.432 0.428 0.432 0.426 0.433 1.043 0.731 0.643 0.571 0.452 0.452 0.594 0.541
720 0.421 0.442 0.425 0.441 0.427 0.445 0.431 0.446 1.104 0.763 0.874 0.679 0.462 0.468 0.831 0.657
Avg 0.375 0.403 0.387 0.405 0.383 0.407 0.387 0.407 0.942 0.684 0.611 0.550 0.414 0.427 0.559 0.515
ECL 96 0.138 0.238 0.144 0.239 0.148 0.240 0.195 0.285 0.219 0.314 0.237 0.329 0.168 0.272 0.197 0.282
192 0.161 0.259 0.160 0.252 0.162 0.253 0.199 0.289 0.231 0.322 0.236 0.330 0.184 0.289 0.196 0.285
336 0.172 0.272 0.176 0.271 0.178 0.269 0.215 0.305 0.246 0.337 0.249 0.344 0.198 0.300 0.209 0.301
720 0.204 0.301 0.207 0.300 0.225 0.317 0.256 0.337 0.280 0.363 0.284 0.373 0.220 0.320 0.245 0.333
Avg 0.169 0.268 0.172 0.266 0.178 0.270 0.216 0.304 0.244 0.334 0.251 0.344 0.192 0.295 0.212 0.300
Exchange 96 0.085 0.204 0.087 0.209 0.086 0.206 0.088 0.205 0.256 0.367 0.094 0.218 0.107 0.234 0.088 0.218
192 0.176 0.300 0.182 0.304 0.177 0.299 0.176 0.299 0.470 0.509 0.184 0.307 0.226 0.344 0.176 0.315
336 0.332 0.418 0.336 0.421 0.331 0.417 0.301 0.397 1.268 0.883 0.349 0.431 0.367 0.448 0.313 0.427
720 0.843 0.691 0.857 0.698 0.847 0.691 0.901 0.714 1.767 1.068 0.852 0.698 0.964 0.746 0.839 0.695
Avg 0.359 0.403 0.366 0.408 0.360 0.403 0.367 0.404 0.940 0.707 0.370 0.413 0.416 0.443 0.354 0.414
Traffic 96 0.425 0.279 0.387 0.262 0.395 0.268 0.544 0.359 0.522 0.290 0.805 0.493 0.593 0.321 0.650 0.396
192 0.455 0.284 0.410 0.270 0.417 0.276 0.540 0.354 0.530 0.293 0.756 0.474 0.617 0.336 0.598 0.370
336 0.467 0.292 0.425 0.277 0.433 0.283 0.551 0.358 0.558 0.305 0.762 0.477 0.629 0.336 0.605 0.373
720 0.505 0.311 0.459 0.296 0.467 0.302 0.586 0.375 0.589 0.328 0.719 0.449 0.640 0.350 0.645 0.394
Avg 0.463 0.292 0.420 0.276 0.428 0.282 0.555 0.362 0.550 0.304 0.760 0.473 0.620 0.336 0.625 0.383
Weather 96 0.155 0.204 0.172 0.214 0.174 0.214 0.177 0.218 0.158 0.230 0.202 0.261 0.172 0.220 0.196 0.255
192 0.208 0.252 0.221 0.259 0.221 0.254 0.225 0.259 0.206 0.277 0.242 0.298 0.219 0.261 0.237 0.296
336 0.267 0.296 0.280 0.300 0.278 0.296 0.278 0.297 0.272 0.335 0.287 0.335 0.280 0.306 0.283 0.335
720 0.344 0.347 0.358 0.349 0.358 0.349 0.354 0.348 0.398 0.418 0.351 0.386 0.365 0.359 0.345 0.381
Avg 0.244 0.275 0.258 0.281 0.258 0.279 0.259 0.281 0.259 0.315 0.271 0.320 0.259 0.287 0.265 0.317
Solar-Energy 96 0.200 0.231 0.208 0.241 0.203 0.237 0.234 0.286 0.310 0.331 0.312 0.399 0.250 0.292 0.290 0.378
192 0.218 0.244 0.232 0.259 0.233 0.261 0.267 0.310 0.734 0.725 0.339 0.416 0.296 0.318 0.320 0.398
336 0.241 0.272 0.262 0.281 0.248 0.273 0.290 0.315 0.750 0.735 0.368 0.430 0.319 0.330 0.353 0.415
720 0.245 0.279 0.258 0.284 0.249 0.275 0.289 0.317 0.769 0.765 0.370 0.425 0.338 0.337 0.356 0.413
Avg 0.226 0.257 0.240 0.266 0.233 0.262 0.270 0.307 0.641 0.639 0.347 0.417 0.301 0.319 0.330 0.401
4.2 Ablations
We conducted detailed ablations on each constituent design of Poly-Mamba as is shown in Table 2. The baseline Mamba4TS [21] for comparison uses the Mamba model with patches as tokens. The results prove that combining LCM and MOPA are important for learning the CDT pattern and neither of them can be omitted. The effectiveness of Order Combining is proved by comparison with Gating only on the outputs of MOPA and LCM (represented as G(LCM,MOPA)) without keeping the low-order trend.
Table 2: Ablations on Poly-Mamba. G(LCM,MOPA) represents Gating only on the outputs of MOPA and LCM, w/o LCM represents only using MOPA and w/o MOPA represents only using LCM.
Models Poly-Mamba G(LCM,MOPA) w/o LCM w/o MOPA Mamba4TS
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh2 96 0.294 0.348 0.298 0.348 0.296 0.348 0.301 0.348 0.297 0.347
192 0.373 0.392 0.383 0.398 0.371 0.391 0.380 0.395 0.392 0.409
336 0.412 0.429 0.419 0.429 0.418 0.429 0.422 0.431 0.424 0.436
720 0.421 0.442 0.428 0.443 0.423 0.442 0.428 0.445 0.431 0.448
Avg 0.375 0.403 0.382 0.405 0.377 0.403 0.383 0.405 0.386 0.410
Exchange 96 0.085 0.204 0.084 0.204 0.088 0.206 0.087 0.206 0.086 0.205
192 0.176 0.300 0.179 0.302 0.185 0.305 0.182 0.304 0.173 0.297
336 0.332 0.418 0.331 0.418 0.334 0.419 0.348 0.428 0.340 0.423
720 0.843 0.691 0.859 0.700 0.855 0.697 0.890 0.711 0.855 0.696
Avg 0.359 0.403 0.363 0.406 0.366 0.407 0.377 0.412 0.364 0.405
Weather 96 0.155 0.204 0.163 0.210 0.164 0.211 0.164 0.210 0.175 0.215
192 0.208 0.252 0.211 0.257 0.209 0.253 0.212 0.256 0.223 0.257
336 0.267 0.296 0.271 0.299 0.269 0.298 0.267 0.296 0.278 0.297
720 0.344 0.347 0.350 0.354 0.347 0.351 0.350 0.350 0.355 0.349
Avg 0.244 0.275 0.249 0.280 0.247 0.278 0.248 0.278 0.258 0.280
ECL 96 0.138 0.238 0.141 0.239 0.156 0.248 0.159 0.258 0.189 0.281
192 0.161 0.259 0.163 0.259 0.168 0.264 0.177 0.275 0.194 0.286
336 0.172 0.272 0.174 0.274 0.178 0.281 0.185 0.284 0.210 0.302
720 0.204 0.301 0.207 0.303 0.210 0.308 0.209 0.304 0.251 0.333
Avg 0.169 0.268 0.171 0.269 0.178 0.275 0.183 0.280 0.211 0.301
4.3 Visual Analysis
Figure 2: (left) One segment of normalized input of Weather dataset. (right) LCM weight matrix L, which is averaged on all encoder layers.
We carry out visual analysis on MOPA and LCM respectively. We visualize the LCM matrix L, which represents the simple linear dependence among channels. For example, as shown in Figure 2, the linear relationship between channels ‘rain’ and ‘raining’ in Weather dataset is more significant than their dependencies on other channels, which is fully in line with the real-world situation.
Figure 3: (left) The first normalized test input of ETTh2 dataset. (middle) The corresponding hidden state value, averaged on all encoder layers, before MOPA operation. (right) The corresponding hidden state value after MOPA operation.
Although MOPA simplifies the complete polynomial expansion and mapping operations, we can observe the coefficients of different channels before and after the MOPA operation to analyze its effect. As is shown in Figure 3, by comparing with the input curve, it can be found that through the mapping of orthogonal polynomial coefficients by MOPA, the coefficients of relevant variables change corresponding to their roles in the fitting. For example, in ETTh2 dataset, MOPA relatively reduces the absolute values of the coefficients of low order terms and relatively increases the absolute values of the coefficients of high order terms for channels ‘MULL’ and ‘LUFL’. This indicates that MOPA finds out more complex CDT patterns than linear relationships between them, and thus more details, fluctuations or nonlinear relationships in the data are captured. As for the channel ‘OT’, whose coefficients of all orders were previously in an averaged state, after MOPA, its relationship with other channels becomes clear and may contain mixed terms, eliminating the weight of its own higher order terms for fitting as before.
4.4 Efficiency
Table 3: Model efficiency comparison. All models are under the same setting. The number of Encoder layers is set to 3, batch size is set to 32, embedding dimension is set to 512, and the dimension of the output linear projection is set to 512. All the experiments are implemented in PyTorch [22] and conducted on a single NVIDIA TELSA V100 32GB GPU.
Dataset Weather (21 Channels)
Model Poly-Mamba S-Mamba Mamba4TS iTransformer Crossformer PatchTST DLinear
Memory(GiB) 0.97 0.64 0.73 0.61 1.18 1.09 0.83
Speed(ms/iter) 27 34 23 21 110 31 28
MSE 0.155 0.172 0.175 0.174 0.158 0.174 0.196
MAE 0.204 0.214 0.215 0.214 0.230 0.218 0.255
Dataset ECL (321 Channels)
Model Poly-Mamba S-Mamba Mamba4TS iTransformer Crossformer PatchTST DLinear
Memory(GiB) 1.43 1.60 0.46 2.31 29.37 11.40 0.46
Speed(ms/iter) 677 108 425 108 2333 507 14
MSE 0.138 0.144 0.189 0.148 0.219 0.195 0.197
MAE 0.238 0.239 0.281 0.240 0.314 0.285 0.282
Similar to Mamba, we cannot use the convolution form in the early SSM (S4 [1]) for fast training. The approach in Mamba [15], which does not actually materialize the full hidden state $h$, is also applicable to Poly-Mamba. Compared with the efficient implementation in Mamba, Poly-Mamba has only newly added the parameter matrices L, M, and scalar Gating weights P. Together with $(\Delta,A,B,C)$, they are directly transferred from slow HBM to fast SRAM, perform the discretization, additional transformation for CDT learning and recurrence in SRAM, and then write the final outputs back to HBM. The work-efficient parallel scan algorithm of Mamba is also applied for avoiding sequential recurrence. Therefore, as is shown in Table 3, the complexity of Poly-Mamba only slightly increases compared to Mamba. However, for the MTSF task, the prediction performance is significantly improved.
Figure 4: The forecasting of the first test time series from the ECL dataset.
5 Conclusion
We delved into the derivation of SSM and found that it has the ability to depict the channel dependency variations with time. For Multivariate Time Series (MTS), we transformed the idea of the original SSM’s orthogonal function basis for real-time approximation of continuously updated functions into real-time approximation of multivariate functions with continuously updated inter-channel dependencies. Based on this, we proposed the Poly-Mamba model, which contains three modules: MOPA for mapping in multivariate orthogonal polynomial basis space, LCM for simple linear mapping between channels, and Order Combining for retaining trends and adaptive fitting. Experimentally, Poly–Mamba achieves state-of-the-art MTS forecasting performance especially when facing MTS scenarios with a large number of channels and real-time changing inter-correlations. In the future, we will continue to further explore the application of our method in other MTS scenarios and explore large-scale time-series pre-training.